Please find out the f(x) = mx + b, given the slope and y-intercept of the linear function
x-intercept -8, y-intercept -1
1
Expert's answer
2014-09-05T14:42:20-0400
Let f(x) = mx + b, It is known that y-intercept=-1, it means that m*0+b=-1, therefore b=-1. It is known that
x-intercept=-8, it means that m*(-8)+b=0, -8m+b=0, i.e. -8m-1=0, -8m=1, whence m=-1/8. Answer: the slope m=-1/8, the y-intercept is b=-1.